one pound of water in changing to steam receives 970.3 BTU at 212°F. Determine the change in specific entropy for the process, BTU/lb-°R
1
Expert's answer
2020-03-13T11:14:32-0400

Energy change =970.3 BTU

Specific entropy = "\dfrac{energy\ change }{mass\times temperature }=\dfrac{970.3}{1\times 212}=4.57 BTU/lb-^oF"